Suppose the risk-free rate is 3.46% and an analyst assumes a market risk premium of 6.46%. Firm A just paid a dividend of $1.29 per share. The analyst estimates the β of Firm A to be 1.25 and estimates the dividend growth rate to be 4.64% forever. Firm A has 288.00 million shares outstanding. Firm B just paid a dividend of $1.89 per share. The analyst estimates the β of Firm B to be 0.78 and believes that dividends will grow at 2.38% forever. Firm B has 182.00 million shares outstanding. What is the value of Firm B?
Answer #1

Risk free rate = 3.46%

market risk premium = 6.46%

Beta of firm B = 0.78

thus, required return(k) of Firm B would be

kb= risk free rate + beta*market risk premium

ke=0.0346 + 0.78*0.0646

kp= 0.084988

Last dividend of Firm B (D0) = $1.89

Constant growth rate of firm B (g) = 2.38%

Current Price of Firm B share would be:

DO *(1 +9) PB = kB - 9

1.89 + (1+0.0238) 0.084988 -0.0238

1.934982 В — о 061188

PB = 31.6235536379682

Outstanding share of Firm B = 182,000,000

Thus,

Value of Firm B would be:

VB = Outstanding Shares * Per share price(PB)

VB = 182,000,000 * 31.6235536379682

Vo = $5,755, 486.762.11

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.